OpenStudy (anonymous):

ln(y) = -2x + c Just did separation of variables. Solving for y (general solution). I cannot for the life of me figure out how to undo a natural log.

OpenStudy (freckles):

e^(ln(y))=y for y>0

OpenStudy (freckles):

\[y=e^{-2x+c}\] you can make answer look prettier

OpenStudy (anonymous):

Woah. e^lny = y? Man, I need to relearn natural log stuff. That's not connecting for me at all

OpenStudy (freckles):

yes f(x)=ln(x) has inverse function g(x)=e^x ans so f((g(x))=x and g(f((x))=x on their domains

OpenStudy (freckles):

\[y=ke^{-2x} \text{ since } e^c \text{ is a constant }\]

OpenStudy (freckles):

is what I meant about making the answer look prettier

OpenStudy (anonymous):

*Whistles* Thanks

OpenStudy (freckles):

lol you whistled at a problem